# Modified Langevin noise formalism for multiple quantum emitters in dispersive electromagnetic environments

### v1

@@ D.1 Finite-dimensional representation
whose support is $S_{p}=T_{p}^{+}\cup T_{p}^{-}$. Its surface divergence is constant over each triangle. Thus, because of the continuity equation, the RWG basis function $\mathbf{f}_{p}$ is associated on the triangle $T_{p}^{\pm}$ to a total electric charge $Q_{p}^{\pm}=\pm\ell_{p}/(i\omega)$, which can be thought as localized in its centroid $\mathbf{g}_{p}^{+}$, and a total magnetic charge $Q_{m}^{\pm}=\pm\ell_{p}/(i\omega\mu_{0})$ localized in its centroid $\mathbf{g}_{p}^{-}$.
@@
$\langle\mathbf{u}|\mathbf{v}\rangle=\int_{\partial\Omega}d^{2}\mathbf{r}\mathbf{u}^{*}(\mathbf{r})\cdot\mathbf{v}(\mathbf{r}).$ (119)
@@ D.3 Average Powers radiated to infinity
To compute the $\mathscr{P}_{ij}^{rad}$ we solve the SIE scattering problem for two excitations: an impressed dipole $\mathbf{p}_{i}$ at $\mathbf{r}_{i}$ and, separately, $\mathbf{p}_{j}$ at $\mathbf{r}_{j}$. Denote the corresponding RWG expansion coefficients by $\{\alpha_{s}^{(i)},\beta_{s}^{(i)}\}$ and $\{\alpha_{t}^{(j)},\beta_{t}^{(j)}\}$. Then, a brute-force but expensive procedure would evaluate the far-field surface integral in Eq. (101) by forming the radiation from the equivalent surface currents at all quadrature points on $S_{\infty}$ for both $\mathbf{p}_{j}$ and $\mathbf{p}_{j}$ excitations, which is an expensive procedure.
@@
with $i=1,\ldots,N_{e}$. For easy of reference, set $\mathbf{p}_{0}^{(i)}=\mathbf{p}_{i}$ and $\mathbf{c}_{0}=\mathbf{r}_{i}$ (There is no impressed magnetic dipole).
@@
$\mathcal{P}_{em}^{ij}(\mathbf{p}_{s},\mathbf{m}_{t})=i\frac{1}{c}\frac{k_{\omega}^{4}}{8\pi\varepsilon_{0}^{2}\zeta_{0}}\\ \mathbf{p}_{s}\times\mathbf{m}^{*}_{t}\cdot\hat{\mathbf{r}}_{st}\,h(k_{\omega}r_{st})$ (124)
@@ Appendix E Lorentz Reciprocity and a far-field identity (alternative proof)
$\mathbf{E}_{i}(\mathbf{r}_{\infty})$ $=\frac{e^{ik_{\omega}r_{\infty}}}{r_{\infty}}\mathbf{A}_{i}(\mathbf{n})+\mathcal{O}(r_{\infty}^{-2})\,$ (126a)
$\mathbf{H}_{i}(\mathbf{r}_{\infty})$ $=\frac{e^{ik_{\omega}r_{\infty}}}{r_{\infty}}\frac{\mathbf{n}\times\mathbf{A}_{i}(\mathbf{n})}{\zeta_{0}}+\mathcal{O}(r_{\infty}^{-2})$ (126b)
@@
$\mathbf{E}_{2}^{inc}$ $=\exp\left(ik_{\omega}\mathbf{n}_{\infty}\cdot\mathbf{r}\right)\mathbf{u}_{\infty}$ (127a)
$\mathbf{H}_{2}^{inc}$ $=\frac{1}{\zeta_{0}}\mathbf{n}_{\infty}\times\mathbf{u}_{\infty}\exp\left(ik_{\omega}\mathbf{n}_{\infty}\cdot\mathbf{r}\right)$ (127b)
@@
$\mathbf{E}_{2}=\mathbf{E}_{2}^{inc}+\mathbf{E}_{2}^{sca};\qquad\mathbf{H}_{2}=\mathbf{H}_{2}^{inc}+\mathbf{H}_{2}^{sca}$ (128)
@@
$\oint_{S_{\infty}}d^{2}\mathbf{r}\,\left(\mathbf{E}_{i}\times\mathbf{H}_{2}-\mathbf{E}_{2}\times\mathbf{H}_{i}\right)\cdot\mathbf{n}=\int_{\mathbb{R}^{3}}d^{3}\;\mathbf{r}\mathbf{E}_{2}\cdot\mathbf{J}_{i}$ (130)
@@
$\frac{1}{\zeta_{0}}\frac{e^{ik_{\omega}r_{\infty}}}{r_{\infty}}\,\oint_{S_{\infty}}d^{2}\mathbf{r}\,\exp\left(ik_{\omega}r_{\infty}\mathbf{n}_{\infty}\cdot\mathbf{n}\right)\\ \left(\mathbf{A}_{i}(\mathbf{n})\times\mathbf{n}_{\infty}\times\mathbf{u}_{\infty}\cdot\mathbf{n}-\mathbf{u}_{\infty}\times\mathbf{n}\times\mathbf{A}_{i}(\mathbf{n})\cdot\mathbf{n}\right)$ (131)
@@
$\frac{1}{\zeta_{0}}{e^{ik_{\omega}r_{\infty}}}\,\int_{0}^{2\pi}d\phi\int_{0}^{\pi}d\theta\,{r_{\infty}}\sin\theta\,\exp\left(ik_{\omega}r_{\infty}\mathbf{n}_{\infty}\cdot\mathbf{n}\right)\\ \left[\mathbf{A}_{i}(\mathbf{n})\times\left(\mathbf{n}_{\infty}\times\mathbf{u}_{\infty}\right)\cdot\mathbf{n}+\mathbf{u}_{\infty}\cdot\mathbf{A}_{i}(\mathbf{n})\right]$ (132)
@@
$\frac{1}{\zeta_{0}}{e^{ik_{\omega}r_{\infty}}}\,\int_{0}^{2\pi}d\phi\int_{0}^{\pi}d\theta\,{r_{\infty}}\sin\theta\,\exp\left(ik_{\omega}r_{\infty}\cos\theta\right)\\ \left[\mathbf{A}_{i}(\theta,\phi)\cdot\mathbf{x}(1+\cos\theta)-\mathbf{A}_{i}(\theta,\phi)\cdot\mathbf{z}\sin\theta\cos\phi\right]$ (133)
The expression 133 contains integrals of the type
@@
$\frac{1}{\zeta_{0}}r_{\infty}{e^{ik_{\omega}r_{\infty}}}{2\pi}\frac{2e^{ik_{\omega}r_{\infty}}}{ik_{\omega}r_{\infty}}\left.\mathbf{A}_{i}\cdot\mathbf{x}\right|_{\theta=0}\\ =-i\frac{4\pi}{\zeta_{0}}\frac{1}{k_{\omega}}e^{2ik_{\omega}r_{\infty}}\left.\mathbf{A}_{i}\cdot\mathbf{x}\right|_{\theta=0}$ (134)
@@
$\left.\mathbf{A}_{i}\cdot\mathbf{x}\right|_{\theta=0}=\frac{k_{\omega}^{2}}{4\pi\varepsilon_{0}}\mu_{i}\,\mathbf{F}_{\omega\mathbf{z}\mathbf{x}}(\mathbf{r}_{i})\cdot\mathbf{u}_{i}$ (135)
@@
Relation  (137) states that the $\mathbf{u}_{\infty}$-polarized far-field amplitude in direction $\mathbf{n}$ is proportional to the projection of the normalized scattering mode $\mathbf{E}_{\omega\mathbf{n}\bm{\nu}}$ at the dipole location onto the dipole orientation $\mathbf{u}_{i}$. In other words, it links the local coupling between the dipole and the scattering mode at $\mathbf{r}_{i}$ to the corresponding far-field pattern produced when the object is driven by a dipole $\mu_{i}\mathbf{u}_{i}$ at $\mathbf{r}_{i}$.

### v2

@@ D.1 Finite-dimensional representation
whose support is $S_{p}=T_{p}^{+}\cup T_{p}^{-}$. Its surface divergence is constant over each triangle. Therefore, by the continuity equation, the RWG basis function $\mathbf{f}_{p}$ is associated with the total electric charge $Q_{p}^{\pm}=\pm\ell_{p}/(i\omega)$ on $T_{p}^{\pm}$, which can be regarded as localized at the corresponding centroid $\mathbf{g}_{p}^{\pm}$, and with the total magnetic charge $Q_{m}^{\pm}=\pm\ell_{p}/(i\omega\mu_{0})$, likewise localized at $\mathbf{g}_{p}^{\pm}$.
@@
$\langle\mathbf{u}|\mathbf{v}\rangle=\int_{\partial V}d^{2}\mathbf{r}\mathbf{u}^{*}(\mathbf{r})\cdot\mathbf{v}(\mathbf{r}).$ (117)
@@ D.3 Average Powers radiated to infinity
To compute the $\mathscr{P}_{ij}^{rad}$ we solve the SIE scattering problem for two excitations: an impressed dipole $\mathbf{p}_{i}$ at $\mathbf{r}_{i}$ and, separately, $\mathbf{p}_{j}$ at $\mathbf{r}_{j}$. Denote the corresponding RWG expansion coefficients by $\{\alpha_{s}^{(i)},\beta_{s}^{(i)}\}$ and $\{\alpha_{t}^{(j)},\beta_{t}^{(j)}\}$. Then, a brute-force but expensive procedure would evaluate the far-field surface integral in Eq. (99) by forming the radiation from the equivalent surface currents at all quadrature points on $S_{\infty}$ for both $\mathbf{p}_{i}$ and $\mathbf{p}_{j}$ excitations, which is an expensive procedure.
@@
with $p=1,\ldots,N_{e}$. For easy of reference, set $\mathbf{p}_{0}^{(i)}=\mathbf{p}_{i}$ and $\mathbf{c}_{0}^{(i)}=\mathbf{r}_{i}$ (There is no impressed magnetic dipole).
@@
$\mathcal{P}_{em}^{ij}(\mathbf{p}_{s},\mathbf{m}_{t})=i\frac{1}{c}\frac{k_{\omega}^{4}}{8\pi\varepsilon_{0}^{2}\zeta_{0}}\mathbf{p}_{s}\times\mathbf{m}^{*}_{t}\cdot\hat{\mathbf{r}}_{st}\,h(k_{\omega}r_{st})$ (122)
@@ Appendix E Lorentz Reciprocity and a far-field identity (alternative proof)
$\mathbf{E}_{i}(\mathbf{r}_{\infty})$ $=\frac{e^{ik_{\omega}r_{\infty}}}{r_{\infty}}\mathbf{A}_{i}(\mathbf{n})+\mathcal{O}(r_{\infty}^{-2}),\,$ (124a)
$\mathbf{H}_{i}(\mathbf{r}_{\infty})$ $=\frac{e^{ik_{\omega}r_{\infty}}}{r_{\infty}}\frac{\mathbf{n}\times\mathbf{A}_{i}(\mathbf{n})}{\zeta_{0}}+\mathcal{O}(r_{\infty}^{-2}),$ (124b)
@@
$\mathbf{E}_{2}^{inc}$ $=\exp\left(ik_{\omega}\mathbf{n}_{\infty}\cdot\mathbf{r}\right)\mathbf{u}_{\infty},$ (125a)
$\mathbf{H}_{2}^{inc}$ $=\frac{1}{\zeta_{0}}\mathbf{n}_{\infty}\times\mathbf{u}_{\infty}\exp\left(ik_{\omega}\mathbf{n}_{\infty}\cdot\mathbf{r}\right),$ (125b)
@@
$\mathbf{E}_{2}=\mathbf{E}_{2}^{inc}+\mathbf{E}_{2}^{sca},\qquad\mathbf{H}_{2}=\mathbf{H}_{2}^{inc}+\mathbf{H}_{2}^{sca},$ (126)
@@
$\oint_{S_{\infty}}d^{2}\mathbf{r}\,\left(\mathbf{E}_{i}\times\mathbf{H}_{2}-\mathbf{E}_{2}\times\mathbf{H}_{i}\right)\cdot\mathbf{n}=\int_{\mathbb{R}^{3}}d^{3}\;\mathbf{r}\mathbf{E}_{2}\cdot\mathbf{J}_{i},$ (128)
@@
$\frac{1}{\zeta_{0}}\frac{e^{ik_{\omega}r_{\infty}}}{r_{\infty}}\,\oint_{S_{\infty}}d^{2}\mathbf{r}\,\exp\left(ik_{\omega}r_{\infty}\mathbf{n}_{\infty}\cdot\mathbf{n}\right)\\ \cdot\left[\mathbf{A}_{i}(\mathbf{n})\times\mathbf{n}_{\infty}\times\mathbf{u}_{\infty}\cdot\mathbf{n}-\mathbf{u}_{\infty}\times\mathbf{n}\times\mathbf{A}_{i}(\mathbf{n})\cdot\mathbf{n}\right].$ (129)
@@
$\frac{1}{\zeta_{0}}{e^{ik_{\omega}r_{\infty}}}\,\int_{0}^{2\pi}d\phi\int_{0}^{\pi}d\theta\,{r_{\infty}}\sin\theta\,\exp\left(ik_{\omega}r_{\infty}\mathbf{n}_{\infty}\cdot\mathbf{n}\right)\\ \cdot\left[\mathbf{A}_{i}(\mathbf{n})\times\left(\mathbf{n}_{\infty}\times\mathbf{u}_{\infty}\right)\cdot\mathbf{n}+\mathbf{u}_{\infty}\cdot\mathbf{A}_{i}(\mathbf{n})\right].$ (130)
@@
$\frac{1}{\zeta_{0}}{e^{ik_{\omega}r_{\infty}}}\,\int_{0}^{2\pi}d\phi\int_{0}^{\pi}d\theta\,{r_{\infty}}\sin\theta\,\exp\left(ik_{\omega}r_{\infty}\cos\theta\right)\\ \cdot\left[\mathbf{A}_{i}(\theta,\phi)\cdot\mathbf{x}(1+\cos\theta)-\mathbf{A}_{i}(\theta,\phi)\cdot\mathbf{z}\sin\theta\cos\phi\right].$ (131)
The expression 131 contains integrals of the type
@@
$\frac{1}{\zeta_{0}}r_{\infty}{e^{ik_{\omega}r_{\infty}}}{2\pi}\frac{2e^{ik_{\omega}r_{\infty}}}{ik_{\omega}r_{\infty}}\left.\mathbf{A}_{i}\cdot\mathbf{x}\right|_{\theta=0}\\ =-i\frac{4\pi}{\zeta_{0}}\frac{1}{k_{\omega}}e^{2ik_{\omega}r_{\infty}}\left.\mathbf{A}_{i}\cdot\mathbf{x}\right|_{\theta=0}.$ (132)
@@
$\left.\mathbf{A}_{i}\cdot\mathbf{x}\right|_{\theta=0}=\frac{k_{\omega}^{2}}{4\pi\varepsilon_{0}}\mu_{i}\,\mathbf{F}_{\omega\mathbf{z}\mathbf{x}}(\mathbf{r}_{i})\cdot\mathbf{u}_{i}.$ (133)
@@
Relation  (135) states that the $\mathbf{u}_{\infty}$-polarized far-field amplitude in direction $\mathbf{n}$ is proportional to the projection of the normalized scattering mode $\mathbf{E}_{\omega\mathbf{n}\boldsymbol{\nu}}$ at the dipole location onto the dipole orientation $\mathbf{u}_{i}$. In other words, it links the local coupling between the dipole and the scattering mode at $\mathbf{r}_{i}$ to the corresponding far-field pattern produced when the object is driven by a dipole $\mu_{i}\mathbf{u}_{i}$ at $\mathbf{r}_{i}$.